A rectangular plate 12 ft long and 8 ft wide is submerged vertically with the longer edge in the surface of the water. Find the force on one side of the plate.


1
Expert's answer
2021-05-03T10:32:52-0400

"F=pS,"

"p=\rho g h,"

"S=hb,"

"F=\rho g h^2b,"

"1~ft=0.3048~m,"

"F=1000\cdot 10\cdot 8^2\cdot 12\cdot 0.3048^3=213~kN."